# Same Content, Different Representations: A Controlled Study for Table QA

### v1

@@ Abstract
Table Question Answering (Table QA) in real-world settings must operate over both structured databases and semi-structured tables containing textual fields. However, existing benchmarks are tied to fixed data formats and have not systematically examined how representation itself affects model performance. We present the first controlled study that isolates the role of table representation by holding content constant while varying structure. Using a verbalization pipeline, we generate paired structured and semi-structured tables, enabling direct comparisons across modeling paradigms. To support detailed analysis, we introduce a diagnostic benchmark with splits along table size, join requirements, query complexity, and schema quality. Our experiments reveal consistent trade-offs: SQL-based methods achieve high accuracy on structured inputs but degrade on semi-structured data, LLMs exhibit flexibility but reduced precision, and hybrid approaches strike a balance, particularly under noisy schemas. These effects intensify with larger tables and more complex queries. Ultimately, no single method excels across all conditions, and we highlight the central role of representation in shaping Table QA performance. Our findings provide actionable insights for model selection and design, paving the way for more robust hybrid approaches suited for diverse real-world data formats.
@@ 2.2 Core Dimensions of Reasoning Difficulty
Table size. Large tables strain token windows and increase attention cost quadratically, leading to truncation, sparse supervision, and difficulty locating relevant rows Liu et al. (2023); Hsieh et al. (2024). LLMs also approximate numbers and struggle with exact comparisons across many values, causing errors in factual and financial QA Chen et al. (2019); Zhu et al. (2021). In this work, we categorize short tables as under 100 rows and long tables as over 100 rows.
@@
3 Fine-grained Diagnostic Benchmark
To systematically study the impact of table representations and disentangle the effects of the four factors outlined in Section 2, we construct a fine-grained diagnostic benchmark. Our design follows two guiding principles: (1) Representation pairing. Comparing structured and semi-structured Table QA requires identical semantics; otherwise, differences may reflect content mismatches. We achieve this by transforming structured tables into semi-structured ones via a controlled verbalization strategy. (2) Factor isolation. We partition existing benchmarks into controlled subsets that vary along the core factors. This allows us to assess each factor’s impact independently.
@@ 4.3 Evaluation
Previous work (Deng et al., 2022; Wu et al., 2025a; Zhu et al., 2021) has largely relied on traditional metrics such as Exact Match (EM) and Partial Match (PM). However, these metrics suffer from well-known limitations: they depend on strict string-level comparison with the gold answer. Consequently, they penalize semantically correct predictions due to paraphrasing, formatting, or minor rounding differences. To address these limitations, we adopt LLMs as judges, a strategy shown to correlate strongly with human evaluation in a variety of reasoning tasks (Maekawa et al., 2025). Concretely, GPT-4o is given both the gold answer and the model’s prediction and prompted to determine semantic correctness. This evaluator tolerates surface variation while enforcing strict semantic equivalence. Unless stated otherwise, all results reported follow this evaluation protocol.
@@ 5.1 RQ1: Model Performance across Paradigms
NL2SQL models achieve the highest accuracy on structured tables, while hybrid models perform best on semi-structured tables. Table 2 shows that input representation has a substantial and consistent effect on performance, with semi-structured inputs posing greater challenges across the methods. On structured inputs, SQL-based methods outperform both hybrids and LLMs, reflecting their reliance on clean, explicit schemas. On semi-structured inputs, this ranking reverses: hybrids lead, LLMs degrade modestly. SQL methods suffer dramatic drops since irregular or implicit structures disrupt symbolic execution.
@@ 5.6 Case Study
We now present illustrative examples that capture the main failure modes when shifting from structured to semi-structured representations. Figure 7 shows a representative case from the “supplier” table, where the query asks for the top ten suppliers by account balance. On the structured version, NL2SQL methods (LLM-NL2SQL and XiYan) execute correctly by leveraging explicit fields such as “s_suppkey” and “s_acctbal”. On the semi-structured version, however, these attributes are embedded into free-text descriptions. The same models either output incorrect results or raise execution errors due to missing column references, illustrating the brittleness of NL2SQL pipelines under representational shifts. A similar failure is observed for the hybrid method Weaver, which also relies on SQL execution and cannot recover when key attributes are verbalized into text. This example highlights the challenges that semi-structured representations pose for models that depend on explicit schema linking and executable queries. We also provide additional case studies in Appendix D that highlight scenarios where different model families diverge in behavior.
@@ Appendix B Effect of Column Selection Settings
Appendix C Fine-grained Results
@@
Table 4: Structured vs. Semi-Structured accuracy across all splits (S1–S5 from BIRD/TableEval, M1–M2 from MMQA). $\Delta$ ($p$) = semi-structured minus structured. Positive = improvement, negative = drop.
@@
Appendix D Additional Case Studies
@@
D.1 Model Behavior on Long Tables
Figure 8 shows a long user profile table with hundreds of rows and a query “What is the id of the youngest user?”. LLM-NL2SQL correctly answers the query on the structured version, which exposes the Age column and allows SQL execution to identify the minimum value (id = 7828). In contrast, LLMs (e.g., GPT-4o) fail to reliably scan the full column and return arbitrary large ids (e.g., 35942). Hybrid methods behave similarly. Weaver returns an arbitrary large id (e.g., 10215), while H-STAR concludes that the task is impossible because of numerous NaN values. This example illustrates that while NL2SQL with execution remains competitive on structured long tables, both LLMs and current hybrid methods are highly sensitive to table length and may produce spurious results when the input context becomes too large.
@@ D.1 Model Behavior on Long Tables
D.2 Model Behavior on Incomplete Schemas
@@
Appendix E Use of Large Language Models

### v2

@@ Abstract
Table Question Answering (Table QA) in real-world settings must operate over both structured databases and semi-structured tables containing textual fields. However, existing benchmarks are tied to fixed data formats and have not systematically examined how representation itself affects model performance. We present the first controlled study that isolates the role of table representation by holding content constant while varying structure. Using a verbalization pipeline, we generate paired structured and semi-structured tables, enabling direct comparisons across modeling paradigms. To support detailed analysis, we introduce RePairTQA, a diagnostic benchmark with splits along table size, join requirements, query complexity, and schema quality. Our experiments reveal consistent trade-offs: SQL-based methods achieve high accuracy on structured inputs but degrade on semi-structured data, LLMs exhibit flexibility but reduced precision, and hybrid approaches strike a balance, particularly under noisy schemas. These effects intensify with larger tables and more complex queries. Ultimately, no single method excels across all conditions, and we highlight the central role of representation in shaping Table QA performance. Our findings provide actionable insights for model selection and design, paving the way for more robust hybrid approaches suited for diverse real-world data formats.
@@ 2.2 Core Dimensions of Reasoning Difficulty
Table size. Large tables strain token windows and increase attention cost quadratically, leading to truncation, sparse supervision, and difficulty locating relevant rows Liu et al. (2023); Hsieh et al. (2024). LLMs also approximate numbers and struggle with exact comparisons across many values, causing errors in factual and financial QA Chen et al. (2019); Zhu et al. (2021). In this work, we categorize short tables as under 100 rows and long tables as over 100 rows. We focus on tables that still fit into the context windows of current LLMs. Extending the framework to ultra-long tables with hundreds or thousands of rows, which likely require chunking or retrieval, is an interesting direction for future work.
@@
3 RePairTQA: Fine-grained Diagnostic Benchmark
To systematically study the impact of table representations and disentangle the effects of the four factors outlined in Section 2, we construct RePairTQA, a fine-grained diagnostic benchmark. Our design follows two guiding principles: (1) Representation pairing. Comparing structured and semi-structured Table QA requires identical semantics; otherwise, differences may reflect content mismatches. We achieve this by transforming structured tables into semi-structured ones via a controlled verbalization strategy. (2) Factor isolation. We partition existing benchmarks into controlled subsets that vary along the core factors. This allows us to assess each factor’s impact independently.
@@ 4.3 Evaluation
Previous work (Deng et al., 2022; Wu et al., 2025a; Zhu et al., 2021) has largely relied on traditional metrics such as Exact Match (EM) and Partial Match (PM). However, these metrics suffer from well-known limitations: they depend on strict string-level comparison with the gold answer. Consequently, they penalize semantically correct predictions due to paraphrasing, formatting, or minor rounding differences. To address these limitations, we adopt LLMs as judges, a strategy shown to correlate strongly with human evaluation in a variety of reasoning tasks (Maekawa et al., 2025). Concretely, GPT-4o is given both the gold answer and the model’s prediction and prompted to determine semantic correctness. This evaluator tolerates surface variation while enforcing strict semantic equivalence. To assess the reliability of our LLM-based evaluator, we randomly sampled 100 evaluation cases across all datasets and model families. Each case was independently annotated by a human and compared to the LLM’s judgment. We observed a 96% agreement rate, demonstrating that the LLM-based evaluation is sufficiently stable for large-scale benchmarking. Unless stated otherwise, all results reported follow this evaluation protocol.
@@ 5.1 RQ1: Model Performance across Paradigms
NL2SQL models achieve the highest accuracy on structured tables, while hybrid models perform best on semi-structured tables. Table 2 shows that input representation has a substantial and consistent effect on performance, with semi-structured inputs posing greater challenges across the methods. On structured inputs, SQL-based methods outperform both hybrids and LLMs, reflecting their reliance on clean, explicit schemas. On semi-structured inputs, this ranking reverses. Hybrid models achieve the highest absolute accuracy on semi-structured tables, while LLMs exhibit relatively modest performance drops. SQL methods suffer dramatic drops since irregular or implicit structures disrupt symbolic execution.
We also examine the robustness of these trends across model scales. For instance, Gemini-2.5-Pro shows the same structured vs. semi-structured performance gap as GPT-4o and other baselines, indicating that sensitivity to representation persists even in state-of-the-art LLMs (full results in Appendix D).
@@ 5.2 RQ2: Impact of Table Size
We also examine the robustness of these trends across table width (Appendix C). We found that reasoning difficulty grows with table width: wide tables (>20 columns) consistently challenge all model families as it makes schema linking harder and context retrieval noisier.
@@ 5.6 Case Study
We now present illustrative examples that capture the main failure modes when shifting from structured to semi-structured representations. Figure 7 shows a representative case from the “supplier” table, where the query asks for the top ten suppliers by account balance. On the structured version, NL2SQL methods (LLM-NL2SQL and XiYan) execute correctly by leveraging explicit fields such as “s_suppkey” and “s_acctbal”. On the semi-structured version, however, these attributes are embedded into free-text descriptions. The same models either output incorrect results or raise execution errors due to missing column references, illustrating the brittleness of NL2SQL pipelines under representational shifts. A similar failure is observed for the hybrid method Weaver, which also relies on SQL execution and cannot recover when key attributes are verbalized into text. This example highlights the challenges that semi-structured representations pose for models that depend on explicit schema linking and executable queries. We also provide additional case studies in Appendix F that highlight scenarios where different model families diverge in behavior.
@@
6 Discussion & Future Work
We summarize the optimal method selection strategy as a decision tree in Figure 8. On the structured data, schema quality is the key discriminator. On clean, descriptive schemas, NL2SQL systems perform best. However, when schemas are incomplete or ambiguous, hybrid pipelines are preferable since their retrieval, normalization, and schema–query alignment steps compensate for missing metadata. On the semi-structured data, the first branch is table size. For long tables, NL2SQL becomes competitive once a stable schema mapping is established, whereas for short tables, query complexity dominates. Simple lookups favor LLM-only reasoning, while compositional or multi-step queries benefit from hybrid pipelines.
A deeper theoretical understanding of representation sensitivity is an important avenue for future work. Our focus here is on establishing a controlled benchmark and documenting empirical trends. Fully explaining the underlying mechanisms will require tools such as architectural ablations, targeted probing of intermediate representations, and instrumentation of model-internal decision pathways. Developing such analyses on top of our benchmark is a promising next step.
Figure 8: Practical decision tree for method selection. NL2SQL excels in structured, clean tables, while Hybrid methods offer superior robustness for semi-structured, noisy, or large-scale contexts. LLMs serve as cost-effective baselines for simple lookups.
@@ Appendix B Effect of Column Selection Settings
Appendix C Effect of Table Width
While our primary analysis differentiates tables by row count, the number of columns (table width) is another important complexity dimension, especially in domains like medical or financial research where high-dimensional data is common. To explore this, we further divide the BIRD diagnostic splits (S1, S3, S4, S5) into three categories based on column count: Narrow (1–5 columns), Medium (6–20 columns), and Wide (>20 columns). Table 4 summarizes the distribution of instances across these categories. Table 5 reports the performance across different paradigms. Broadly, we observe an inverse correlation between table width and accuracy across most methods.
On structured tables, performance degrades consistently as width increases. For instance, GPT-4o drops from 52.75% on narrow tables to 27.91% on wide tables. SQL-based methods demonstrate superior resilience on wide tables, likely because symbolic SQL queries can effectively select relevant columns. On semi-structured tables, the challenge is amplified for LLMs. Models like GPT-4o and H-STAR suffer significant drops on wide tables. Interestingly, NL2SQL maintains consistent performance, further validating that symbolic execution provides a stable fallback when the input representation becomes noisy or high-dimensional.
Table 4: Distribution of table width buckets across BIRD-based diagnostic splits.
Width Bucket Column Range # Instances
Narrow 151-5 455
Medium 6206-20 329
Wide >20>20 89
Table 5: Performance comparison across table width buckets. Bold indicates the best performance within each setting and bucket.
Structured Accuracy (%) Semi-Structured Accuracy (%)
Model Narrow Medium Wide Narrow Medium Wide
GPT-4o 52.75 37.80 27.91 48.79 37.92 25.58
Gemini-2.5-flash 51.87 43.47 31.46 54.29 41.34 31.46
Qwen3-235B 42.64 34.35 28.09 43.96 31.91 30.34
LLM-NL2SQL 76.48 65.55 57.30 48.79 43.90 47.19
XiYan 75.16 64.94 56.18 24.18 35.06 43.82
H-STAR 43.30 32.01 25.84 44.84 32.52 21.35
Weaver 54.41 44.07 38.20 57.71 51.06 33.71
Appendix D Additional Results on a Large Instruction-Tuned LLM: Gemini-2.5-Pro
To validate the generalizability of our findings, we evaluated Gemini-2.5-Pro using the same setup as in Section 3.2. The results are summarized in Table 6. While Gemini-2.5-Pro demonstrates strong overall capabilities, achieving high accuracy on simpler structured queries (e.g., 92.63% on S1 Structured), it shows the same sensitivity to representation as GPT-4o and other baselines. We found a consistent accuracy drop when moving from structured to semi-structured tables across all difficulty levels. For instance, on the most complex subset (S5), the accuracy significantly declines from 41.31% (Structured) to 33.98% (Semi-Structured).
Table 6: Performance of Gemini-2.5-Pro across all splits (S1-S5 and M1-M2). The model consistently shows performance degradation on semi-structured inputs.
Split Structured Acc. (%) Semi-Structured Acc. (%) $\Delta$ ($p$)
S1 92.63 88.42 -4.21
S2 81.07 73.37 -7.70
S3 76.59 73.86 -2.73
S4 60.76 59.49 -1.27
S5 41.31 33.98 -7.33
M1 82.44 78.45 -3.99
M2 35.54 29.52 -6.02
Appendix E Fine-grained Results
@@
Table 7: Structured vs. Semi-Structured accuracy across all splits. $\Delta$ ($p$) = semi-structured minus structured.
@@
Appendix F Additional Case Studies
@@
F.1 Model Behavior on Long Tables
Figure 9 shows a long user profile table with hundreds of rows and a query “What is the id of the youngest user?”. LLM-NL2SQL correctly answers the query on the structured version, which exposes the Age column and allows SQL execution to identify the minimum value (id = 7828). In contrast, LLMs fail to reliably scan the full column and return arbitrary large ids. Hybrid methods behave similarly. Weaver returns an arbitrary large id, while H-STAR concludes that the task is impossible because of numerous NaN values. This example illustrates that while NL2SQL with execution remains competitive on structured long tables, both LLMs and current hybrid methods are highly sensitive to table length and may produce spurious results when the input context becomes too large.
@@ F.1 Model Behavior on Long Tables
F.2 Model Behavior on Incomplete Schemas
@@
Appendix G Use of Large Language Models